# Nearly-zero large-angle anisotropy of the cosmic microwave background

### v5

@@ Abstract
The global isotropy of the universe is analyzed on the scale of the cosmic horizon, using the angular correlation function $C(\Theta)$ of cosmic microwave background (CMB) temperature at large angular separation $\Theta$. Even-parity correlation $C_{\rm even}(\Theta)$ is introduced to obtain a direct, precise measure of horizon-scale curvature anisotropy independent of the unknown dipole, with uncertainty dominated by models of Galactic emission. In maps from Planck, $C_{\rm even}(\Theta)$ at $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ is found to be much closer to zero than previously documented measurements of correlation, with an absolute value three to four orders of magnitude smaller than expected in standard theory. Such a small variation from zero is estimated to occur by chance in a fraction $\simeq 10^{-4.3}$ to $\simeq 10^{-2.8}$ of standard realizations. We consider an alternative interpretation of this result, as a signature of a scale-invariant angular symmetry of cosmological initial conditions. An example of such a symmetry is formulated geometrically, by assuming that quantum fluctuations during inflation have spacelike coherence bounded by compact causal diamonds, and become classical perturbations when world lines cross inflationary horizons. It is argued that this interpretation is inconsistent with the standard theory of initial conditions, but is broadly consistent with other cosmological measurements, and is subject to further tests.
@@
pubyear: 2025pagerange: Nearly-zero large-angle anisotropy of the cosmic microwave backgroundLABEL:lastpage
@@ 1 Introduction
Both the near-perfect isotropy and the smaller-scale departures from uniformity that lead to large scale cosmic structure can in principle be generated by a causal physical process if there is an early inflationary acceleration of the cosmological scale factor, which allows any two locations to have causal contact at a sufficiently early time. In standard inflation theory, departures from uniformity arise from quantum fluctuations. The standard quantum model of these fluctuations (Baumann, 2011; Weinberg, 2008), based on local effective quantum field theory (QFT), agrees with measurements of cosmic structure over a wide range of scales, including CMB correlations on scales less than a few degrees.
@@
However, this theory does not agree very well with the precise large-scale uniformity measured on larger angular scales. Indeed, it has been realized since the earliest measurements of CMB anisotropy, with the COBE satellite (Bennett et al., 1994; Hinshaw et al., 1996), that the universe is actually much smoother at large $\Theta$ than inflation theory typically predicts. The unexpectedly small magnitude of $C(\Theta)$ at large angular separation was confirmed with higher precision in subsequent studies with WMAP and Planck (Bennett et al., 2003, 2011; Planck Collaboration, 2016, 2020b).
@@
In standard cosmology, the small correlation is attributed to a statistical fluke of our particular realized sky; that is, only a small fraction of realizations that agree with structure on smaller scales are as smooth as the real sky on the largest scales (Copi et al., 2009; Muir et al., 2018). Even though large-angle anisotropy provides the most direct probe of initial conditions, $C(\Theta)$ at $\Theta$ larger than a few degrees, or equivalently the angular power spectrum $C_{\ell}$ at angular wavenumber $\ell$ less than about thirty, are generally disregarded in tests of cosmological models. This practice makes sense within the standard framework of inflationary theory, which predicts many possible realizations of the CMB sky that differ significantly from each other on large angular scales.
@@
In spite of a theoretical expectation that casts doubt on its significance, the structure of CMB anisotropy on the largest scales nevertheless remains a unique phenomenon, which preserves a precisely measurable pattern of cosmic initial conditions from the earliest time and largest distance accessible to us. With this special status as the primary motivation, the current paper aims to measure correlations on large scales, using all-sky CMB temperature maps, up to current limits of systematic measurement uncertainty.
At present, these limits are mainly determined by the best comprehensive all-sky models of Galactic foregrounds, developed by the Planck team. Previous studies have not reached this measurement limit, due to another fundamental systematic uncertainty: because CMB maps have unknown dipole ($\ell=1$) harmonic modes subtracted, their measured $C(\Theta)$ differs from that of the CMB horizon itself (Peebles, 2022).
@@
In this paper, we study the even-parity part of the angular correlation function $C_{\rm even}(\Theta)$, which gives a direct estimate of true horizon-scale correlations, independent of the unknown dipole or any other model parameter, over a wider range of angles. Our new measurements of $C_{\rm even}(\Theta)$ are plotted for several maps in Fig. (1), together with 100 random realizations of the standard model. From the plot it can be immediately seen that the absolute value of $C_{\rm even}(\Theta)$ over a range of angles near $\Theta\simeq 90^{\circ}$ is much smaller than expected, and differs from zero no more than the different maps differ from each other. Quantitatively, the surprising new result is that the variance of $C_{\rm even}(\Theta)$ from zero in the angular range $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ is three to four orders of magnitude smaller than expected in the standard cosmological model.
@@
Over the range $\Theta\simeq 90^{\circ}\pm 15^{\circ}$, the measured isotropy is not just surprising in the context of standard theory, it is extraordinary in absolute terms: to illustrate with one example from the estimates below, the fractional dimensionless correlation residual of the smoothest Planck map (NILC) is $\int C_{\rm even}^{2}/(2.7{\rm K})^{4}\sim 2\times 10^{-26}$. In other physical systems, such very small dimensionless numbers can often be traced to the deep mathematical structure of the system. We are bound to ask whether the measured near-perfect isotropy apparently preserved in the CMB could possibly signify a fundamental symmetry of cosmic initial conditions that is not included in standard theory.
A fundamental symmetry that accounts for vanishing angular correlation would have to be a property of any sky, and any realization of initial conditions. It also needs to account for the specific range of angular scales, $\Theta\simeq 90^{\circ}\pm 15^{\circ}$, where nearly-vanishing even-parity correlation is observed. We show here that it is geometrically possible to formulate such a symmetry, defined by covariant relativistic geometrical relationships, that could account for large-angle $C(\Theta)$ measurements, and at the same time agree with the nearly scale-invariant 3D power spectrum of cosmological perturbations on all scales.
Our formulation is based on a scale-invariant principle of causal coherence widely tested in entangled laboratory quantum systems (Zeilinger, 1999; Vilasini & Renner, 2024): namely, that no correlation can occur between systems contained within completely separate regions of space-time, because they are causally independent. A quantum process that occurs entirely in the future of one event, and entirely in the past of some future event at the same location, can modify physical relationships only within a unique 4D region of space-time bounded by their light cones, called a causal diamond. Quantum processes contained within completely separate causal diamonds do not produce physical correlations with each other. A scale- and conformally-invariant symmetry of angular correlation in cosmic initial conditions could arise from physical potential differences generated on inflationary horizons from quantum fluctuations completely contained within causal diamonds during cosmic inflation.
@@
The angular symmetry is formulated here from the standard conformal geometry of classical relativistic cosmology, but it is not compatible with the part of inflationary theory that depends on the standard QFT model for quantum fluctuations and initial perturbations. It should be noted in this regard that the standard formulation of initial conditions in the QFT model has previously been criticized on theoretical grounds (Penrose, 1989; Ellis, 1999; Ijjas & Steinhardt, 2016): for example, an initially perfectly uniform and very large comoving patch of a perfectly uniform 3D background space is simply assumed as an initial condition very early during inflation, and specification of a particular initial vacuum state on this spacelike surface completely determines the final pattern of realized perturbations. Moreover, at a basic physical level, the standard QFT model assumes that field modes evolve coherently over the entire initial patch throughout inflation, even at spacelike separations much larger than the inflationary horizon, in spite of known inconsistencies in QFT on scales larger than causal horizons (Cohen et al., 1999; Hollands & Wald, 2004; Stamp, 2015). If large-angle cosmic correlations preserve unique information about gravitational quantum states on horizons, the exceptional symmetry of CMB isotropy may be more physically profound than it appears to be in the familiar context of QFT-based inflation theory (Hogan, 2019).
The plan of this paper is as follows. In Sec. 2 we describe a new measurement of large-angle even-parity CMB correlations. In Sec. 3, we review the relativistic causal structure of inflationary cosmology, and present a covariant formulation of causal relationships that could define symmetries of angular correlation. In Sec. 4 we compare measurements with standard QFT realizations at angular separations where causal symmetries could lead to zero correlation. In Sec. 5, we summarize the alternative interpretations of the data. An overall summary is presented in Sec. 6. In the Appendix (Sec. 7), we describe in more detail how inflation with causally-coherent symmetry differs from the standard QFT picture, and could modify standard concordance cosmology.
@@ 2.1 Angular spectrum and correlation function
The power spectrum $C_{\ell}$ is the statistical tool generally used for tests of cosmological models. However, $C(\Theta)$ provides a more direct signature of geometrical causal relics of initial conditions. Because it correlates azimuthal averages at a polar angle $\Theta$ (Eq. 6), sharp boundaries of causal correlation within spherical horizon surfaces produce exactly zero $C(\Theta)$ in a range of $\Theta$ determined by the polar angle of their circular intersection. Causal boundaries are not apparent in $C_{\ell}$, even though it contains the same statistical information about the angular distribution.
@@ 2.2 Scalar perturbations and large-scale CMB anisotropy
On large angular scales, temperature anisotropy in the CMB is largely determined by primordial scalar curvature perturbations $\Phi$ of the cosmological metric on a thin sphere at the location of the last scattering surface (Sachs & Wolfe, 1967). Apart from the Doppler-induced dipolar anisotropy from local motion, their angular distributions on scales larger than a few degrees are the same:
@@
In this sense, CMB correlation provides a direct measurement of angular symmetries in initial conditions.
Gravity also introduces anisotropy on large angular scales much later during the dark energy dominated era, the late-time integrated Sachs-Wolfe effect (ISW) (Hu & Dodelson, 2002; Francis & Peacock, 2010). This effect is generated by primordial perturbations in 3D, as the CMB light propagates through space on our past light cone. In the linear regime, it is also determined by the invariant local scalar potential $\Phi$ that preserves its original primordial spatial distribution from the end of inflation. However, anisotropy from this effect comes from perturbations at comoving distances much smaller than the last scattering surface, and from correlations of matter at different radii. Thus, the measured total CMB anisotropy depends on angular cross-correlations of $\Phi$ in three comoving spatial dimensions.
In the interpretation below, we will neglect other physical effects, such as radiation transport and Doppler motion at recombination, which do not modify the angular spectrum significantly at spherical harmonics with $\ell\lesssim 30$ (Hu & Dodelson, 2002).
@@ 2.3 Data
As explained in Hagimoto et al. (2020), we use all-sky CMB maps made with subtracted models of Galactic emission, in order to minimize correlation artifacts introduced by masks. Our analysis is based on foreground-corrected maps of the CMB temperature based on the fifth and third public release databases of the WMAP and Planck collaborations, respectively. For WMAP, we use the ILC map with and without the fitted monopole included. In the case of Planck, we use several different maps based on different techniques for modeling the Galaxy. Recognizing that the noise properties of the foreground-corrected maps are not well characterized and that the 2-point function is correlated between angles, we use the variation between foreground subtraction methods and experiments as a proxy for correlation function uncertainty. We only compare integrated residuals of measured values and standard model realizations of 2-point correlation functions.
@@ 2.4 Measured correlation function
In particular, the measured variation of $C_{\rm even}(\Theta)$ from zero in the range $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ is orders of magnitude smaller than previously documented correlations. Its absolute value is comparable with differences between different foreground-subtracted maps, that is, apparently as close to zero as current measurements allow. For the three Planck maps— which arguably provide the most accurate models of Galactic foregrounds— the correlation is nearly indistinguishable from zero on the scale plotted in Fig. (1).
@@
Figure 2: Spacetime diagram illustrating causal relationships betweeen events on world lines $A$ and $B$ spaced a comoving distance $\eta_{0}$ apart. Causal correlations can form where the future light cone of an initial event intersects the past light cone of an observer. In the figure, these relationships are causal diamonds, examples of which are the blue and green diamonds. While the relationships in the figure hold in general for any two world lines, to analyze the situation of the observation of the CMB, we imagine the observer to be on world line $A$ at a time $+\eta_{0}$ and set the time $\eta=0$ to be the end of inflation and shortly after, at the recombination epoch. Thus $\eta_{0}$ is the current horizon distance and the CMB is emitted by a 3-sphere with radius $\eta_{0}$, marked by the red ellipse in the figure and denoted $\mathcal{S}_{A}(\eta_{0})$. The goal is to analyze what events during inflation can, even in principle, generate correlations in the CMB and determine the pattern of the correlations. Because of the causal constraint, not all correlations generated by the standard model for CMB fluctuations are present in this picture. CMB temperature anisotropy on large angular scales is dominated by gravitational perturbations near the last scattering surface, near ${\cal S}_{A}(\eta_{0})$, but also has contributions from distortions arising from later perturbations along the light cone at $\eta>0$, shown as ISW in the figure. Incoming information to the world lines $A$ and $B$ during inflation is bounded by inflationary horizons ${\cal H}_{A}$ and ${\cal H}_{B}$. Shaded 4D regions represent causal diamonds bounded by the comoving 2-spheres ${\cal S}_{A}(\eta_{0}/2)$ (blue) and ${\cal S}_{B}(\eta_{0})$ (green). We posit that coherent quantum fluctuations are bounded by causal diamonds and convert into classical potential differences on inflationary horizons, so a coherent perturbation of ${\cal S}_{A}(\eta_{0}/2)$ forms within the blue causal diamond shown, which starts on $A$ at $-\eta_{0}$; this is the last causal diamond whose fluctuations entangle $A$ with $B$, or with any other point at $|\vec{r}-\vec{r}_{A}|=\eta_{0}$. At locations $\vec{r}$ with radius $|\vec{r}-\vec{r}_{A}|<\eta_{0}/2$, coherent fluctuations generate correlations of scalar potential $\Phi(\vec{r})$ with $\Phi(\vec{r}_{A})$ which are independent of correlations of $\Phi(\vec{r}_{A})$ with $\Phi(\vec{r}_{B})$, so angular correlations vanish in a range of $\Theta$ around $\pi/2$, as shown in Fig. (3).

### v7

# Dipole-independent measurements of nearly-zero CMB correlation: a possible symmetry of primordial causal quantum coherence

@@ Abstract
Anisotropy of space-time is measured on the scale of the cosmic horizon, using the angular correlation function $C(\Theta)$ of cosmic microwave background (CMB) temperature at large angular separation $\Theta$. Even-parity correlation $C_{\rm even}(\Theta)$ is introduced to obtain a direct, precise measure of horizon-scale curvature anisotropy independent of the unknown dipole, with uncertainty dominated by models of Galactic emission. In maps from WMAP and Planck, $C_{\rm even}(\Theta)$ at $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ is found to be much closer to zero than in previously documented measurements. Variation from zero as small as that in the Planck maps is estimated to occur by chance in a fraction $\simeq 10^{-4.3}$ to $\simeq 10^{-2.8}$ of standard realizations. Measurements are found to be consistent with zero correlation in a range of angles expected from quantum fluctuations during inflation whose spacelike coherence is bounded by inflationary horizons around every location at every epoch. This scale-invariant symmetry of cosmological initial conditions is incompatible with the standard quantum theory of initial conditions, but is broadly consistent with other cosmological measurements, and is subject to further tests.
@@
pubyear: 2025pagerange: Dipole-independent measurements of nearly-zero CMB correlation: a possible symmetry of primordial causal quantum coherenceLABEL:lastpage
@@ 1 Introduction
Both the near-perfect isotropy and the smaller-scale perturbations that lead to large scale cosmic structure can in principle be generated by a causal physical process if there is an early inflationary acceleration of the cosmological scale factor, which allows any two locations to have causal contact at a sufficiently early time. In standard inflation theory, perturbations arise from quantum fluctuations. The standard model (Baumann, 2011; Weinberg, 2008), based on Gaussian noise derived from local effective quantum field theory (QFT), agrees with measurements of cosmic structure over a wide range of scales, including CMB correlations on scales less than a few degrees.
@@
However, this theory does not agree very well with the precise isotropy measured on larger angular scales. Indeed, it has been realized since the earliest measurements of CMB anisotropy, with the COBE satellite (Bennett et al., 1994; Hinshaw et al., 1996), that the universe is actually much smoother at large $\Theta$ than the Gaussian model typically predicts. The unexpectedly small magnitude of $C(\Theta)$ at large angular separation was confirmed with higher precision in subsequent studies with WMAP and Planck (Spergel et al., 2003; Bennett et al., 2011; Planck Collaboration, 2016, 2020b, 2020c).
@@
Even though large-angle anisotropy provides the most direct probe of initial conditions, correlations at $\Theta$ larger than a few degrees, or equivalently the angular power spectrum $C_{\ell}$ at angular wavenumber $\ell$ less than about thirty, are often disregarded in tests of cosmological models, because the standard Gaussian model predicts many possible realizations of the CMB sky that differ significantly from each other on large angular scales. In standard cosmology, the small large-angle correlation is attributed to a statistical fluke of our particular realized sky; that is, only a small fraction of realizations that agree with structure on smaller scales are as smooth as the real sky on the largest scales. This persistent anomaly has led some authors to question the standard picture (Copi et al., 2007, 2009; Copi et al., 2015; Muir et al., 2018).
@@
In spite of a theoretical expectation that casts doubt on its significance, the structure of CMB anisotropy on the largest scales nevertheless remains a unique phenomenon, which preserves a precisely measurable intact pattern of cosmic initial conditions. Anomalously small large-angle anisotropy may be a direct signature of physics that shaped initial conditions.
Previous theoretical studies have explored possible fundamental implications of anomalous large-angle isotropy. In one notable study, Copi et al. (2019) studied toy models of primordial perturbations with compact support limited by a characteristic length, based on compact wavelets on 3D comoving spacelike hypersurfaces in initial conditions— a real-space modification of the Gaussian noise usually modeled by an initial QFT vacuum state defined in comoving harmonic 3-space. They compare models with CMB maps and suggest that the initial perturbations themselves may have a suppressed correlation function on large length scales, which “could signal that when inflationary perturbations are generated, they are coherent only over distances shorter than the horizon due to inflationary microphysics.”
The measurements reported in this paper are designed to test a related hypothesis about microphysical constraints on inflationary correlations, formulated here as a covariant bound in four dimensions instead of a compact correlation in 3D. It is based on a physical principle of causal coherence, that quantum systems in causally disconnected regions of space-time lead to independent outcomes. In our application of this principle, physical correlations of perturbations from every epoch during inflation are bounded by inflationary horizons around every comoving location. Such a causal bound is a natural consequence of the kind of compact nonlocal causal spacelike coherence familiar in physical systems with “spooky” quantum correlations, but it is incompatible with standard inflationary QFT formulations of gravitational fluctuations that depend on separability of space and time for their modal decomposition. [^1] Causal coherence in 4D is also incompatible with models that impose constraints on initial correlations to be compact in comoving real 3-space, as in the toy wavelet model of Copi et al. (2019). As shown below, scale-invariant causally coherent correlations in 4D lead to nonzero 3D correlations, as well as exotic constraints on angular correlations, at all comoving spacelike separations. We show below how covariant causal coherence in four dimensions can lead to angular correlations from primordial gravitational perturbations that are not only small, but exactly vanish over a specific range of angular separation around $\Theta=90^{\circ}$.
The new measurements reported in this paper are designed to test this angular symmetry. In particular, we introduce separate measurements of even and odd parity correlations to allow direct model-independent tests of exact universal null symmetries of angular correlation that do not occur in QFT models.
The plan of this paper is as follows. In Sec. 2 we further explain the motivation and design of the current study. In Sec. 3 we describe a new measurement of large-angle even-parity CMB correlations. In Sec. 4, we review the relativistic causal structure of inflationary cosmology, and analyze covariant causal bounds in four dimensions that can lead to relict symmetries of angular correlation. In Sec. 5 we compare measurements with standard QFT realizations at angular separations where causal symmetries could lead to zero correlation. In Sec. 5.3, we summarize the alternative interpretations of the data. An overall summary is presented in Sec. 6. In the Appendix (Sec. 7), we describe in more detail how causally-coherent inflation differs from the standard QFT picture, and give some examples of other ways it could modify standard concordance cosmology.
2 Motivation and design of this study
The main goal of the current study is to use CMB correlations on the largest scales to study possible exact symmetries of cosmological initial conditions. This motivation drives specific choices for measurements.
Systematic uncertainty in measurements of large-angle CMB temperature correlations is mainly determined by the effect of Galactic foregrounds. A common practice for precision studies is to mask off large parts of the sky where models of the Galaxy are unreliable. This works well for some statistical tests, and with a fully defined Gaussian model such as standard inflation. However, in general the unmeasured masked sections of sky generate biases that make them unsuitable for null tests (Hagimoto et al., 2020). Instead, we adopt all-sky models of the Galaxy developed by the Planck and WMAP teams using a variety of different approaches, and use the differences between them as a guide to systematic uncertainties.
Another fundamental systematic measurement uncertainty arises because CMB maps have unknown dipole ($\ell=1$) harmonic modes subtracted, so that their measured $C(\Theta)$ differs from that of the CMB horizon itself (Copi et al., 2019; Peebles, 2022). This uncertainty may be addressed in a parameter-free way by separate measurement of the even parity component of correlation $C_{\rm even}(\Theta)$, which has no dipole. This measurement is well suited to test a unique symmetry-based prediction for $C_{\rm even}(\Theta)=0$, where there is no particular theoretical expectation for the dipole.
@@ 2 Motivation and design of this study
In this paper, we extend this study to the whole even-parity part of the angular correlation function $C_{\rm even}(\Theta)$, which gives a direct estimate of true horizon-scale correlations, independent of the unknown dipole or any other model parameter, over a wider range of angles. Our new measurements of $C_{\rm even}(\Theta)$ are plotted for several maps in Fig. (1), together with 100 random realizations of the standard model. From the plot it can be immediately seen that the absolute value of $C_{\rm even}(\Theta)$ over a range of angles near $\Theta\simeq 90^{\circ}$ is still much smaller than expected, and differs from zero no more than the different maps differ from each other. Quantitatively, the surprising new result derived below is that the variance of $C_{\rm even}(\Theta)$ from zero in the angular range $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ is three to four orders of magnitude smaller than expected in the standard cosmological model.
@@
A small dimensionless number in nature can often be traced to a fundamental symmetry. We are bound to ask whether the measured near-perfect isotropy apparently preserved in CMB correlation could possibly signify a fundamental symmetry of initial perturbations that is not preserved in the standard theory.  [^2] Large-scale uniformity ultimately depends on a symmetry of the initial state. The actual measured large-angle uniformity is extraordinary when expressed in absolute terms. To illustrate with one example from estimates below, the fractional dimensionless correlation residual of the smoothest Planck map (NILC) around $\Theta\sim 90^{\circ}$, expressed as a fractional perturbation of total curvature on the scale of the horizon, is $\int C_{\rm even}^{2}/(2.7{\rm K})^{4}\sim 2\times 10^{-26}$. In standard theory, fluctuations are expected to add variance several orders of magnitude larger than this.
A fundamental symmetry that accounts for vanishing angular correlation would have to be a property of any sky, and any realization of initial conditions. It also needs to account for the specific range of angular scales, $\Theta\simeq 90^{\circ}\pm 15^{\circ}$, where nearly-vanishing even-parity correlation is observed. We show below that it is geometrically possible to formulate such a symmetry, defined by covariant relativistic geometrical relationships, that could account for large-angle $C(\Theta)$ measurements, and at the same time agree with the nearly scale-invariant 3D power spectrum of cosmological perturbations on all scales.
Our formulation is based on a principle of causal coherence widely tested in entangled laboratory quantum systems (Zeilinger, 1999; Vilasini & Renner, 2024): namely, that no correlation can occur between systems contained within completely separate regions of space-time, because they are causally independent. A quantum process that occurs entirely in the future of one event, and entirely in the past of some future event at the same location, can modify physical relationships only within a unique 4D region of space-time bounded by their light cones, called a causal diamond. Quantum processes contained within completely separate causal diamonds do not produce physical correlations with each other. A scale- and conformally-invariant symmetry of angular correlation in cosmic initial conditions could arise from physical potential differences generated on inflationary horizons from quantum fluctuations that are completely contained within causal diamonds during cosmic inflation.
@@
The angular symmetry is formulated here from the standard conformal geometry of classical relativistic cosmology, but it is not compatible with the standard QFT model for quantum fluctuations and initial gravitational perturbations. The standard formulation of initial conditions has previously been challenged on theoretical grounds (Penrose, 1989; Ellis, 1999; Ijjas & Steinhardt, 2016), and it is well known that modifications of QFT are expected in a deeper theory of quantum gravity that addresses entanglement with causal horizons (Cohen et al., 1999; Hollands & Wald, 2004; Stamp, 2015). If large-angle cosmic correlations preserve unique information about nonlocal, causal superpositions of gravitational quantum states on horizons that do not occur in QFT, the exceptional symmetry of CMB isotropy may be more physically profound than it appears to be in the familiar context of QFT-based inflation theory (Hogan, 2019).
@@ 3.1 Angular spectrum and correlation function
The power spectrum $C_{\ell}$ is the statistical tool generally used for tests of cosmological models. However, $C(\Theta)$ provides a more direct signature of geometrical causal relics of initial conditions, for reasons discussed below. Causal boundaries are not apparent in $C_{\ell}$, even though it contains the same statistical information about the angular distribution.
@@ 3.2 Scalar perturbations and large-scale CMB anisotropy
On large angular scales, temperature anisotropy in the CMB is mostly determined by primordial scalar curvature perturbations $\Phi$ of the cosmological metric on a thin sphere at the location of the last scattering surface (Sachs & Wolfe, 1967). Apart from the Doppler-induced dipolar anisotropy from local motion, their angular distributions on scales larger than a few degrees are the same:
@@
In this sense, CMB correlation provides a direct measurement of any angular symmetry of initial conditions on a particular sphere.
Gravity also introduces some anisotropy during propagation, via the integrated Sachs-Wolfe effect (ISW) (Hu & Dodelson, 2002; Francis & Peacock, 2010; Copi et al., 2016). This effect is generated by primordial perturbations in 3D, as the CMB light propagates through space on our past light cone. In the linear regime, it is also determined by the invariant local scalar potential $\Phi$ that preserves its original primordial spatial distribution from the end of inflation. Anisotropy from this effect comes from perturbations at comoving distances smaller than the last scattering surface. Thus, symmetries of CMB temperature correlation are mainly determined by the SW effect of $\Phi$ on the last scattering surface (Eq. 8), with a relatively small additional ISW contribution shaped by angular cross-correlations of $\Phi$ in three comoving spatial dimensions (Copi et al., 2016).
In the analysis below, we will neglect other physical effects, such as radiation transport and Doppler motion at recombination, which do not modify the angular spectrum significantly at spherical harmonics with $\ell\lesssim 30$ (Hu & Dodelson, 2002).
@@ 3.3 Data
As explained in Hagimoto et al. (2020), we use all-sky CMB maps made with subtracted models of Galactic emission, in order to minimize correlation artifacts introduced by masks. Our analysis is based on foreground-corrected maps of the CMB temperature based on the fifth and third public release databases of the WMAP and Planck collaborations, respectively. For WMAP, we use the ILC map with and without the fitted monopole included [^3] Both of these possibilities were presented as models by the WMAP team, so we show them here for completeness. In fact, this monopole must be interpreted as an artifact of imperfect Galaxy model subtraction: there can be no actual measured monopole of true CMB anisotropy by definition, since the monopole is isotropic.. In the case of Planck, we use several different maps based on different techniques for modeling the Galaxy. Recognizing that the noise properties of the foreground-corrected maps are not well characterized and that the 2-point function is correlated between angles, we use the variation between foreground subtraction methods and experiments as a proxy for correlation function uncertainty. We only compare integrated residuals of measured values and standard model realizations of 2-point correlation functions.
@@ 3.4 Measured correlation function
In particular, the measured variation of $C_{\rm even}(\Theta)$ from zero in the range $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ is orders of magnitude smaller than previously documented correlations. Its absolute value is comparable with differences between different foreground-subtracted maps, that is, apparently as close to zero as current measurements allow. All three Planck maps— which arguably provide the most accurate models of Galactic foregrounds— are nearly indistinguishable from zero on the scale plotted in Fig. (1).
@@
Figure 2: Four dimensional spacetime in conformal coordinates, showing causal relationships between world lines $A$ and $B$ spaced a comoving distance $2\eta_{0}$ apart. While the relationships in the figure hold in general for any two world lines and any scale, to analyze the situation of the observation of the CMB, we imagine the observer to be on world line $A$ at a time $\eta_{0}$ and set the time $\eta=0$ to be the end of inflation and shortly after, at the recombination epoch. Thus $\eta_{0}$ is the current horizon distance and the CMB is emitted by a 3-sphere with radius $\eta_{0}$, denoted $\mathcal{S}_{A}(\eta_{0})$. The goal is to analyze what events during inflation can, even in principle, generate correlations in the CMB and determine the pattern of the correlations. CMB temperature anisotropy on large angular scales is dominated by gravitational perturbations near the last scattering surface, near ${\cal S}_{A}(\eta_{0})$, but also has contributions from distortions arising from later perturbations along the light cone at $\eta>0$, shown as ISW in the figure. Incoming information to the world lines $A$ and $B$ during inflation is bounded by inflationary horizons ${\cal H}_{A}$ and ${\cal H}_{B}$. Shaded 4D regions represent causal diamonds bounded by the comoving 2-spheres ${\cal S}_{A}(\eta_{0})$ (blue) and ${\cal S}_{B}(2\eta_{0})$ (green). We posit that coherent quantum fluctuations are bounded by causal diamonds and convert into classical potential differences on inflationary horizons, so a coherent perturbation of ${\cal S}_{A}(\eta_{0})$ forms within the blue causal diamond shown, which starts on AA at $2\eta_{0}$; this is the last causal diamond whose fluctuations entangle AA with $B$. At locations $\vec{r}$ with radius $|\vec{r}-\vec{r}_{A}|<\eta_{0}$, correlations of $\Phi(\vec{r})$ with $\Phi(\vec{r}_{A})$ are independent of correlations of $\Phi(\vec{r}_{A})$ with $\Phi(\vec{r}_{B})$, so the intersection of spheres ${\cal S}_{A}(\eta_{0})$ and ${\cal S}_{B}(2\eta_{0})$ represents the boundary of nonzero correlation measured at $(\vec{r}_{A},\eta_{0})$. A $B$ horizon at larger distance intersects ${\cal S}_{A}(\eta_{0})$ at a larger angle, but its fluctuations are not entangled with ${\cal S}_{A}(\eta_{0})$. This leads to a maximum angular separation for causal correlation, shown in Fig. (3).